OpenStudy (anonymous):

How do I get the term (4/(x-1)) <= x+2 to a single rational expression?

OpenStudy (anonymous):

I bring it all over the the left side, to get \[(4/(x-1)) -x - 2 <= 0\], then I multiple the second part by (x-1)/(x-1) to get (x-2)(x-1)/(x-1)...then??

OpenStudy (unklerhaukus):

instead of multiplying by \(\frac{x-1}{x-1}\) multiply by \(\frac{x+1}{x+1}\)

OpenStudy (anonymous):

\[4/(x-1) - (x ^{2} -x - 2)/(x+1) \]?

OpenStudy (unklerhaukus):

\[\frac 4{x-1} \leq x+2\]\[\frac 4{x-1} - x+2\leq0\] \[\left(\frac 4{x-1}\right)\times\frac{x+1}{x+1} - (x+2)\times\frac{x+1}{x+1}\leq0\] \[\frac {4({x+1})}{(x-1)\times(x+1)}- \frac{(x+2)(x+1)}{x+1} \leq0\] \[\frac {4({x+1})}{(x-1)(x+1)}- \frac{(x+2)(x+1)(x-1)}{(x+1)(x-1)} \leq0\] \[\frac {4({x+1})-(x+2)(x+1)(x-1)}{(x+1)(x-1)} \leq0\]

OpenStudy (anonymous):

I think you have taken it seriously @UnkleRhaukus

OpenStudy (anonymous):

Thanks so much for the amazingly clear and helpful answer. You're my hero.

OpenStudy (anonymous):

How did you know what to multiply everything by though?

OpenStudy (unklerhaukus):

when we are trying to get rid of a term like this in the denominator like \((x+a)\), multiply by its conjugate \((x-a)\)

OpenStudy (unklerhaukus):

well ,we really multiply by \(\frac{x-a}{x-a}=1\)

OpenStudy (anonymous):

\[\frac 4{x-1} \leq x+2 \implies \frac{4}{x - 1} -x - 2 \le 0 \implies \frac{4 - (x+2)(x - 1)}{(x-1)} \le 0\] \[\frac{4 - (x^2 + x - 2)}{(x-1)} \le 0 \implies \frac{x^2 + x + 2}{(x-1)} \ge 0\]

OpenStudy (anonymous):

Ha ha ha.. Done something wrong..

OpenStudy (unklerhaukus):

how can we get a rational expression?

OpenStudy (anonymous):

Meaning ??

OpenStudy (unklerhaukus):

my result wasent rational, and neither is yours @waterineyes

OpenStudy (anonymous):

WolframAlpha says that it can be reduced to \[(x-2)(x+3)/(x-1)\] (How do you insert fractions?)

OpenStudy (anonymous):

Sorry I have done there a mistake: \[4 - x^2 - x + 2 \implies -x^2 -x + 6 \implies -(x^2 + x - 6) \implies -(x+3)(x-2)\]

OpenStudy (anonymous):

So that will become: \[\frac{-(x+3)(x-2)}{(x-1)} \le 0 \implies \frac{(x+3)(x-2)}{(x-1)} \ge 0\]

OpenStudy (anonymous):

OH! Haha you make it look easy. Back to the books for me! :)